# On Edge Coloring of Multigraphs

### v4

@@ 5.2.1 A few basic properties
Let $\varphi_{\eta}=\varphi/(G-T_{n+1},\alpha,\eta)$. Note that $P_{2}$ is a $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{2})$. So $f_{i}$ is colored by $\beta$ under both $\varphi$ and $\varphi_{\gamma}$ for $i=1,2$. Consider $\sigma=\varphi_{\eta}/P_{2}$. Note that $\beta\in\overline{\sigma}(b_{2})$, and $f_{1}$ is colored by $\beta$. We claim that
(5) $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\gamma}$, and $T_{n,j}(b_{2})$ is an ETT with $(n-1)$-rungs and a good hierarchy of $(j-1)$ levels under $\sigma$.
Note that $\sigma$ is $(T_{n,j}(b_{2})-b_{2},D_{n},\varphi_{\eta})$-stable, it is an ETT with $(n-1)$-rungs under $\sigma$ by (2), (4) and Corollary 3.1, once we show that $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$. Indeed, because $\beta$ is not closed for $T_{n,j}$ under $\varphi$, we have $\beta\neq\gamma_{h}$ by (a). So by (4), both $\eta$ and $\beta$ are different from $\gamma_{h}$. Recall that we have $\beta\neq\delta_{h}$ in the current case. So if $\eta\neq\delta_{h}$, then $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\eta}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$ by (A2.5). If $\eta=\delta_{h}$, then since $b_{2}\notin T_{n,j-1}$ by (3) and the elementariness of $T_{n,j}$ under $\varphi$, we have $\delta_{h}\notin\overline{\varphi}(T_{n,j})$. Thus $\delta_{h},\gamma_{h}\notin\Gamma^{j}$ by Definition 4. So all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi_{\eta}$ are contained in $G[V(T_{n,j})]$ by Definition 4 (iv), (2) and (4). Thus by the construction of $\sigma$, they are colored the same under $\sigma$. Since $T_{h}$ only has one $(\gamma_{h},\delta_{h})$-exit path under $\varphi_{\eta}$ by (A2), $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\eta}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$ by (A2.5). It is routine to see that $T_{n,j}(b_{2})$ has a good hierarchy of $(j-1)$ levels, see Remark 3.
Recall that $P_{2}$ is an $(\alpha,\beta)$-exit path of $T_{n,j}$ while $\alpha$ is closed in $T_{n,j}$ under $\varphi$, we have
@@
Let $T^{\prime}=T_{n+1}(b_{3})+f_{1}$ under $\sigma$. Because $\beta\notin\Gamma^{j-1}$ by (6), we see that $T^{\prime}$ satisfies Definition 4 (i), and thus $T^{\prime}$ has a good hierarchy of $(j-1)$ levels, see Remark 3. Let $T^{\prime}_{n,j}$ be a tree-sequence obtained from $T^{\prime}$ by the following algorithm:
@@
(7) $T^{\prime}_{n,j}$ is an ETT with a good hierarchy of $(j-1)$-levels under $\sigma$ and $T^{\prime}_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed.
@@
Case II:
$\beta=\delta_{h}$.
@@
(8) $\varphi_{\eta}$ is $T_{n}$ mod $\varphi$, and therefore $T_{n,j}$ is an ETT with a good hierarchy of $(j-1)$-levels under $\varphi_{\eta}$.
Observe that $P_{2}$ is a $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{2})$. So similarly as in the previous case, we let $\sigma=\varphi_{\eta}/P_{2}$. We claim that
@@
Note that following (9), with the same argument as showing (6), we have $\beta\notin\overline{\varphi}(T_{n,j}(b_{2}))$ (so $Q^{\prime}(T_{n,j})=1$ in this case), $\beta\notin\Gamma^{j-1}$, and $T_{n,j}(b_{2})$ is an ETT with a good hierarchy of $(j-1)$-levels, and therefore similarly as earlier, none of $a_{1}$ and $a_{2}$ are in $T^{\prime}_{n,j}$ also an ETT with a good hierarchy of $(j-1)$-levels ($T^{\prime}_{n,j}$ is constructed the same way as in Case I). So $P_{1}$ and $P_{2}$ give us two $(\beta,\eta)$-paths intersecting $T^{\prime}_{n,j}$, and they provides two $(\beta,\eta)$-exit paths $L_{1}$ and $L_{2}$ of $T^{\prime}$, respectively. Moreover, $L_{1}$ provide exit vertex in $T^{\prime}_{n,j}$ after $b_{2}$ along $\prec$ of $T^{\prime}_{n,j}$, as $T_{n,j}(b_{2})+f_{1}\subseteq T^{\prime}_{n,j}$. Since $\delta_{h}=\beta\in\overline{\sigma}(b_{2})$, $(\sigma,T^{\prime}_{n,j},\beta,\eta,L_{1},L_{2})$ with $\beta\neq\gamma_{n}$ provides a counter-example with $Q^{\prime}(T^{\prime}_{n,j})=0$, a contradiction.
Following (a), the proof of (9) is identical to the proof of (7) in Lemma 4.1 with $T_{n,j}$ in place of $T_{n+1}$. Here we give a summary of the actual algorithm used here. Suppose that $T_{h}$ has a different $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ from the $(\gamma_{h},\delta_{h})$-exit vertex $v_{2}$ under $\varphi_{\eta}$, where $\sigma=\varphi_{\eta}/P_{2}$. In the case that $v_{1}\prec v_{2}$, we let $n^{\prime}$ be the smallest index such that $v_{2}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is RE finished. Then under $\varphi_{\gamma}$, we have a contradiction to Lemma 4.4 for $P$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $\gamma$ in place of $\alpha$, $v_{2}$ in place of $x$, $v_{1}$ in place of $y$, and $\gamma_{h},\delta_{h}$ in place of $\gamma,\delta$.
In the case that $v_{2}\prec v_{1}$, we let $n^{\prime}$ be the smallest index such that $v_{1}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is RE finished. Pick $\tau\in\overline{\varphi}(x_{e})$ with $\tau\neq\gamma_{h^{\prime}}$ and $\alpha^{\prime}\in\overline{\varphi}(v_{1})$. Let $\mu=\sigma/(G-T_{h},\gamma,\tau)$. Let $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma_{h})$ if $\tau\neq\gamma_{h}$, and $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma)$ otherwise. Let $\mu_{2}$ be obtained from $\mu_{1}$ by switching all the $\eta$-colored edges of $P$ not in $P_{x_{e}}(\eta,\tau,\mu_{1})$ to $\tau$ by Kempe changes. Then under $\mu_{2}$, we reach a contradiction to Lemma 4.6 with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $D^{\prime}$ in place of $D$, $\alpha^{\prime}$ in place of $\alpha$, $\delta_{h}$ in place of $\delta$, $x_{e}$ in place of $r$, $v_{1}$ in place of $x$, $v_{2}$ in place of $y$, and $b_{2}$ in place of $w$. ∎
@@
Let $\alpha\in\overline{\varphi}(T_{n,j})$ for some $0\leq j\leq q$ with $\alpha\neq\gamma_{n}$, and $\beta$ is any color in $[k]$. Suppose $P$ is an $(\alpha,\beta)$-path not intersecting $T_{n,j}$. If $\pi=\varphi/P$, then $\pi$ is $T_{n}+f_{n}$ mod $\varphi$.
@@
Let $h$ be the initial index for $\Theta_{n}$ and $\pi=\varphi/P$. Then $\gamma_{n}=\gamma_{h}$, $\delta_{n}=\delta_{h}$, and $\pi$ is $(T_{n},D_{n},\varphi)$-stable. We propose to show (1) $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$. Note that by assuming (1), we immediately see that $\pi$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) for $T_{n+1}$ in place of $T_{n+2}$. In the case that $\alpha\neq\delta_{h}$ and $\beta\notin S_{h}=\{\gamma_{h},\delta_{h}\}$, (1) holds trivially. If $\alpha=\delta_{h}$, then all the $(\gamma_{h},\delta_{h})$-chains intersecting $T_{h}$ are contained in $G[V(T_{n})]$ by TAA, and (1) also holds as $\pi$ is $(T_{n},D_{n},\varphi)$-stable. Thus we may assume $\alpha\notin S_{h}$ and $\beta\in S_{h}$. In this case, we may assume one of $\gamma$ and $\tau$, say $\gamma\neq\gamma_{h}$. Moreover, since $\gamma\in\overline{\varphi}(x_{e})$, $\gamma\neq\delta_{h}$. Following Claim 5.1, $\alpha$ is closed for some $T_{n,j^{\prime}}$ with $0\leq j^{\prime}\leq j$. By slightly abuse of notation, we may assume that $\alpha$ is closed for $T_{n,j}$. By Claim 5.2 (a), $T_{n,j}$ is also closed for $\gamma$. So $\sigma=\varphi(G-T_{n,j},\alpha,\gamma)$ is also $(T_{n,j},D_{n},\varphi)$-stable. As $\gamma,\alpha\notin S_{h}$, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as under $\varphi$. So $\sigma$ is also $T_{n}+f_{n}$ mod $\varphi$ by (A2.5). Note that under $\sigma$, $P$ is a $(\gamma,\beta)$-path not intersecting $T_{n,j}$. So by (A3.5) for $T_{n}$, $\pi_{1}=\sigma/P$ is $T_{n}+f_{n}$ mod $\sigma$, and therefore is $T_{n}+f_{n}$ mod $\varphi$ too. Note that $\pi=\pi_{1}/(G-T_{n,j},\alpha,\gamma)$, we similarly see that $\pi$ is $T_{n}+f_{n}$ mod $\pi_{1}$ as $\gamma,\alpha\notin S_{h}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, as desired. ∎
@@
For any $y\in V(T(y_{p-1})-T_{n,q})$, $|\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}-\varphi(T(y)-T_{n,q})|\geq 2+2n$. Furthermore, for $i=1,2,3$, there exist distinct colors $\beta_{i}\in\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}$ with $\beta_{i}\notin\varphi(T(y)-T_{n,q})$ such that either $\beta_{i}\notin\Gamma^{q}\cup D_{n,q}$ or $\beta_{i}\in\Gamma^{q}_{r}$ for some $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$ (here $\delta_{r}$ might be different for different $i$). Consequently, for any $y=y_{k}$ with $20\leq k\leq p-$ 2, there exists a color $\beta\in\overline{\varphi}(T(y))$ with $\beta\notin\varphi(T(y_{k+2})-T_{n,q})$ such that either $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$ or $\beta\in\Gamma^{q}_{r}$ for some $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$.
@@
Since under $\varphi$, $T_{n}$ is elementary following (5.0) and $T_{l}$ is closed for each $l\in[n]$, we have $|T_{l}|$ being odd, and thus $|\overline{\varphi}(T_{n})\setminus\{\gamma_{n}\}|\geq 2n+2$ following Lemma 3.2. So $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+3$ when $q\geq 1$. Because $|V(T(y)-T_{n,q})|=|E(T(y)-T_{n,q})|$, following (5.0), we have $|\overline{\varphi}(T(y)-T_{n,q})|\geq|\varphi(T(y)-T_{n,q})|$. Note that in the case that $q=0$, the connecting edge $f_{n}$ is colored by $\delta_{n}\notin\overline{\varphi}(T_{n})$, so $|\varphi(T(y)-T_{n,q})\cap\overline{\varphi}(T(y))|<|\varphi(T(y)-T_{n,q})|$ unless $\delta_{n}\in\overline{\varphi}(T(y))$. Therefore, $|\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}-\varphi(T(y)-T_{n,q})|\geq|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq|\overline{\varphi}(T_{n})\setminus\{\gamma_{n}\}|\geq 2n+2$, and equality holds only if $q=0$ and $\delta_{n}\in\overline{\varphi}(T(y))$. Now let us show the “furthermore” part. Denote $\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}$ by $C$. Now let $|C-\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q})|=d$. Clearly we have as desired if $d\geq 3$, so we may assume $d<3$. Since
$C-\varphi(T(y)-T_{n,q})=(C-\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q}))\cup((\Gamma^{q}\cup D_{n,q})\cap C-\varphi(T(y)-T_{n,q})\cap C),$
@@
$|(\Gamma^{q}\cup D_{n,q})\cap C-\varphi(T(y)-T_{n,q})\cap C|\geq|C-\varphi(T(y)-T_{n,q})|-|C-\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q})|\geq 2n+2-d,$
and equality holds only if $q=0$ and $\delta_{n}\in\overline{\varphi}(T(y))$. Since $|D_{n,q}|\leq n$ and $d<3$, by the Pigeonhole Principle with $D_{n,q}$ many pigeonholes, there are at least $(3-d)$ distinguished colors $\beta\in\Gamma^{q}_{r}\cap C$ such that $\beta\notin\varphi(T(y)-T_{n,q})$ with $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$ (different $\beta$ can have different $\delta_{r}$). Together with $|C-\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q})|=d$, the “furthermore” part holds. Finally we prove the “consequently” part of the current claim. Since there are only two edges in $T(y_{k+2})-T(y_{k})$, clearly it holds by the “furthermore” part if $k\geq 1$. So we assume $k=0$, that is, $y=y_{0}$ being the last vertex of $T_{n,q}$. Note that if $q\geq 1$, we have $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+3$. Since there are only two edges in $T(y_{2})-T_{n,q}$, $D_{n,q}\notin\overline{\varphi}(T_{n,q})$, and $|\Gamma^{q}|\leq 2n$, we have a color $\beta\in\overline{\varphi}(T_{n,q})$ with $\beta\notin\varphi(T(y_{2})-T_{n,q})$ such that $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$, as desired. So $q=0$, and we have $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+2$ in this case. However, in this case the first edge on $T(y_{2})-T_{n,q}$ is colored by $\delta_{n}\notin\overline{\varphi}(T_{n})$. So again, since there are only two edges in $T(y_{2})-T_{n,q}$, $D_{n,q}\notin\overline{\varphi}(T_{n,q})$, and $|\Gamma^{q}|\leq 2n$, we also have a desired color $\beta\in\overline{\varphi}(T_{n,q})$ with $\beta\notin\varphi(T(y_{2})-T_{n,q})$ such that $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$. ∎

### v5

@@ 5.2.1 A few basic properties
Let $\varphi_{\eta}=\varphi/(G-T_{n+1},\alpha,\eta)$. Note that $P_{2}$ is an $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{2})$, and $f_{i}$ is colored by $\beta$ under $\varphi_{\eta}$ for $i=1,2$. Consider $\sigma=\varphi_{\eta}/P_{2}$. So $\beta\in\overline{\sigma}(b_{2})$, and $f_{1}$ is still colored by $\beta$ under $\sigma$. We claim that
(5) $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$, and $T_{n,j}(b_{2})$ is an ETT with (n1)(n-1)-rungs and a good hierarchy of (j1)(j-1) levels under $\sigma$.
Note that $\sigma$ is $(T_{n,j}(b_{2})-b_{2},D_{n},\varphi_{\eta})$-stable, it is an ETT with (n1)(n-1)-rungs under $\sigma$ by (2), (4) and Corollary 3.1, once we show that $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$. Indeed, because $\beta$ is not closed for $T_{n,j}$ under $\varphi$, we have $\beta\neq\gamma_{h}$ by (a). So by (4), both $\eta$ and $\beta$ are different from $\gamma_{h}$. Recall that we have $\beta\neq\delta_{h}$ in the current case. So if $\eta\neq\delta_{h}$, then $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\eta}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$ by (A2.5). If $\eta=\delta_{h}$, then since $b_{2}\notin T_{n,j-1}$ by (3) and the elementariness of $T_{n,j}$ under $\varphi$, we have $\delta_{h}\notin\overline{\varphi}(T_{n,j})$. Thus $\delta_{h},\gamma_{h}\notin\Gamma^{j}$ by Definition 4. So all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi_{\eta}$ are contained in $G[V(T_{n,j})]$ by Definition 4 (iv), (2) and (4). Thus by the construction of $\sigma$, they are colored the same under $\sigma$. Since $T_{h}$ only has one $(\gamma_{h},\delta_{h})$-exit path under $\varphi_{\eta}$ by (A2), $T_{h}$ still has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as $\varphi_{\eta}$, and thus $\sigma$ is $T_{n}+f_{n}$ mod $\varphi_{\eta}$ by (A2.5). It is routine to see that $T_{n,j}(b_{2})$ has a good hierarchy of (j1)(j-1) levels, see Remark 4.
So $T_{n,j}(b_{2})$ is elementary under $\sigma$ following (5) and (5.0). Thus $\beta\notin\overline{\sigma}(T_{n,j}(b_{2})-b_{2})$ by the construction of $\sigma$. Moreover, we have
@@
Let $T^{\prime}=T_{n+1}(b_{3})+f_{1}$ under $\sigma$. Because $\beta\notin\Gamma^{j-1}$ by (6), we see that $T^{\prime}$ satisfies Definition 4 (i), and thus $T^{\prime}$ has a good hierarchy of (j1)(j-1) levels, see Remark 4. Let $T^{\prime}_{n,j}$ be a tree-sequence obtained from $T^{\prime}$ by the following algorithm:
@@
(7) $T^{\prime}_{n,j}$ is an ETT with a good hierarchy of (j1)(j-1) levels under $\sigma$ and $T^{\prime}_{n,j}$ is $(\cup_{\delta_{m}\in D_{n,j}}\Gamma^{j-1}_{m})^{-}$-closed.
@@
Case II: $\beta=\delta_{h}$.
@@
(8) $\varphi_{\eta}$ is $T_{n}$ mod $\varphi$, and therefore $T_{n,j}$ is an ETT with a good hierarchy of (j1)(j-1) levels under $\varphi_{\eta}$.
Observe that $P_{2}$ is an $(\eta,\beta)$-path under $\varphi_{\eta}$ with $\eta\in\overline{\varphi}_{\eta}(b_{2})$. So similarly as in the previous case, we let $\sigma=\varphi_{\eta}/P_{2}$. We claim that
@@
Note that following (9), with the same argument as showing (6), we have $\beta\notin\overline{\varphi}(T_{n,j}(b_{2}))$ (so $Q^{\prime}(T_{n,j})=1$ in this case), $\beta\notin\Gamma^{j-1}$, and $T_{n,j}(b_{2})$ is an ETT with a good hierarchy of (j1)(j-1) levels, and therefore similarly as earlier, none of $a_{1}$ and $a_{2}$ are in $T^{\prime}_{n,j}$ also an ETT with a good hierarchy of (j1)(j-1) levels ($T^{\prime}_{n,j}$ is constructed the same way as in Case I). So $P_{1}$ and $P_{2}$ give us two $(\beta,\eta)$-paths intersecting $T^{\prime}_{n,j}$, and they provides two $(\beta,\eta)$-exit paths $L_{1}$ and $L_{2}$ of $T^{\prime}$, respectively. Moreover, $L_{1}$ provide exit vertex in $T^{\prime}_{n,j}$ after $b_{2}$ along $\prec$ of $T^{\prime}_{n,j}$, as $T_{n,j}(b_{2})+f_{1}\subseteq T^{\prime}_{n,j}$. Since $\delta_{h}=\beta\in\overline{\sigma}(b_{2})$, $(\sigma,T^{\prime}_{n,j},\beta,\eta,L_{1},L_{2})$ with $\beta\neq\gamma_{n}$ provides a counterexample with $Q^{\prime}(T^{\prime}_{n,j})=0$, a contradiction.
Following (a) and (3), the proof of (9) is almost identical to the proof of (7) in Lemma 4.1 with $T_{n,j}$ in place of $T_{n+1}$. Here we give a summary of the actual algorithm used here. Suppose that $T_{h}$ has a different $(\gamma_{h},\delta_{h})$-exit vertex $v_{1}$ under $\sigma$ from the $(\gamma_{h},\delta_{h})$-exit vertex $v_{2}$ under $\varphi_{\eta}$, where $\sigma=\varphi_{\eta}/P_{2}$. In the case that $v_{1}\prec v_{2}$, we let $n^{\prime}$ be the smallest index such that $v_{2}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is $\operatorname{RE}$ finished. Then under $\varphi_{\gamma}$, we have a contradiction to Lemma 4.4 for $P$ with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $\gamma$ in place of $\alpha$, $v_{2}$ in place of $x$, $v_{1}$ in place of $y$, and $\gamma_{h},\delta_{h}$ in place of $\gamma,\delta$.
In the case that $v_{2}\prec v_{1}$, we let $n^{\prime}$ be the smallest index such that $v_{1}\in T_{n^{\prime}}$ and $T_{n^{\prime}}$ is $\operatorname{RE}$ finished. Let $h^{\prime}$ be the initial index of $T_{n^{\prime}}$. Pick $\alpha^{\prime}\in\overline{\varphi}(v_{1})$. Recall $\gamma\neq\gamma_{h}$ (see the assumption before (3)), following (3), we let $\mu$ be a coloring obtained from $\sigma$ by interchanging labeling for $\gamma$ and $\tau$ if $\gamma\neq\gamma_{h^{\prime}}$, and let $\mu=\sigma/(G-T_{h},\gamma,\tau)$ if $\gamma=\gamma_{h^{\prime}}$. Either way, we have $\tau\neq\gamma_{h^{\prime}}$, and $\sigma(P)\subseteq\{\tau,\eta,\delta_{h}\}$. We then let $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma_{h})$ if $\tau\neq\gamma_{h}$, and $\mu_{1}=\mu/(G-T_{h},\alpha^{\prime},\gamma)$ otherwise. Let $\mu_{2}$ be obtained from $\mu_{1}$ by switching all the $\eta$ colored edges of $P$ not in $P_{x_{e}}(\eta,\tau,\mu_{1})$ to $\tau$ by Kempe changes. Then under $\mu_{2}$, we reach a contradiction to Lemma 4.6 with $T_{n^{\prime}}$ ($n^{\prime}\leq n$) in place of $T_{n+1}$, $\alpha^{\prime}$ in place of $\alpha$, $\delta_{h}$ in place of $\delta$, $x_{e}$ in place of $r$, $v_{1}$ in place of $x$, $v_{2}$ in place of $y$, and $b_{2}$ in place of $w$. ∎
@@
Let $\alpha\in\overline{\varphi}(T_{n,j})$ for some $0\leq j\leq q$ with $\alpha\neq\gamma_{n}$, and $\beta$ be any color in $[k]$. Suppose $P$ is an $(\alpha,\beta)$-path not intersecting $T_{n,j}$. If $\pi=\varphi/P$, then $\pi$ is $T_{n}+f_{n}$ mod $\varphi$.
@@
Let $h$ be the initial index for $\Theta_{n}$ and $\pi=\varphi/P$. Then $\gamma_{n}=\gamma_{h}$, $\delta_{n}=\delta_{h}$, and $\pi$ is $(T_{n},D_{n},\varphi)$-stable. We propose to show
(1) $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$.
Note that by assuming (1), we immediately see that $\pi$ is $T_{n}+f_{n}$ mod $\varphi$ by (A2.5) for $T_{n+1}$ in place of $T_{n+2}$.
In the case that $\alpha\neq\delta_{h}$ and $\beta\notin S_{h}=\{\gamma_{h},\delta_{h}\}$, (1) holds trivially. If $\alpha=\delta_{h}$, then all the $(\gamma_{h},\delta_{h})$-chains intersecting $T_{h}$ are contained in $G[V(T_{n})]$ by TAA, and (1) also holds as $\pi$ is $(T_{n},D_{n},\varphi)$-stable.
Thus we may assume $\alpha\notin S_{h}$ and $\beta\in S_{h}$. In this case, we may assume one of $\gamma$ and $\tau$, say $\gamma\neq\gamma_{h}$. Moreover, since $\gamma\in\overline{\varphi}(x_{e})$, $\gamma\neq\delta_{h}$. Following Claim 5.1, $\alpha$ is closed for some $T_{n,j^{\prime}}$ with $0\leq j^{\prime}\leq j$. By slightly abuse of notation with $j$ denoting $j^{\prime}$, we may assume that $\alpha$ is closed for $T_{n,j}$. By Claim 5.2 (a), $T_{n,j}$ is also closed for $\gamma$. So $\sigma=\varphi(G-T_{n,j},\alpha,\gamma)$ is also $(T_{n,j},D_{n},\varphi)$-stable. As $\gamma,\alpha\notin S_{h}$, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\sigma$ as under $\varphi$. So $\sigma$ is also $T_{n}+f_{n}$ mod $\varphi$ by (A2.5). Note that under $\sigma$, $P$ is a $(\gamma,\beta)$-path not intersecting $T_{n,j}$. So by (A3.5) for $T_{n}$, $\pi_{1}=\sigma/P$ is $T_{n}+f_{n}$ mod $\sigma$, and therefore is $T_{n}+f_{n}$ mod $\varphi$ too. Note that $\pi=\pi_{1}/(G-T_{n,j},\alpha,\gamma)$, we similarly see that $\pi$ is $T_{n}+f_{n}$ mod $\pi_{1}$ as $\gamma,\alpha\notin S_{h}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$, as desired.
@@
For any $y\in V(T(y_{p-1})-T_{n,q})$, $|(\overline{\varphi}(T(y))\setminus\{\gamma_{n}\})\setminus\varphi(T(y)-T_{n,q}))|\geq 2+2n$. Furthermore, for $i=1,2,3$, there exist distinct colors $\beta_{i}\in\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}$ with $\beta_{i}\notin\varphi(T(y)-T_{n,q})$ such that either $\beta_{i}\notin\Gamma^{q}\cup D_{n,q}$ or $\beta_{i}\in\Gamma^{q}_{r}$ for some $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$ (here $\delta_{r}$ might be different for different $i$). Consequently, for any $y=y_{k}$ with $0\leq k\leq p-2$, there exists a color $\beta\in\overline{\varphi}(T(y))$ with $\beta\notin\varphi(T(y_{k+2})-T_{n,q})$ such that either $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$ or $\beta\in\Gamma^{q}_{r}$ for some $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$.
@@
Since under $\varphi$, $T_{n}$ is elementary following (5.0) and $T_{l}$ is closed for each $l\in[n]$, we have $|T_{l}|$ being odd, and thus $|\overline{\varphi}(T_{n})\setminus\{\gamma_{n}\}|\geq 2n+2$ following Lemma 3.2. So $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+3$ when $q\geq 1$. Because $|V(T(y)-T_{n,q})|=|E(T(y)-T_{n,q})|$, following (5.0), we have $|\overline{\varphi}(T(y)-T_{n,q})|\geq|\varphi(T(y)-T_{n,q})|$. Note that in the case that $q=0$, the connecting edge $f_{n}$ is colored by $\delta_{n}\notin\overline{\varphi}(T_{n})$, so $|\varphi(T(y)-T_{n,q})\cap\overline{\varphi}(T(y))|<|\varphi(T(y)-T_{n,q})|$ unless $\delta_{n}\in\overline{\varphi}(T(y))$. Therefore, $|(\overline{\varphi}(T(y))\setminus\{\gamma_{n}\})\setminus\varphi(T(y)-T_{n,q})|\geq|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq|\overline{\varphi}(T_{n})\setminus\{\gamma_{n}\}|\geq 2n+2$, and equality holds only if $q=0$ and $\delta_{n}\in\overline{\varphi}(T(y))$.
Now let us show the “furthermore” part. Denote $\overline{\varphi}(T(y))\setminus\{\gamma_{n}\}$ by $C$. Now let $|C\setminus(\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q}))|=d$. Clearly we have as desired if $d\geq 3$, so we may assume $d<3$. Since
$C\setminus\varphi(T(y)-T_{n,q})=(C\setminus(\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q})))\cup(((\Gamma^{q}\cup D_{n,q})\cap C)\setminus(\varphi(T(y)-T_{n,q})\cap C)),$
@@
$|((\Gamma^{q}\cup D_{n,q})\cap C)\setminus(\varphi(T(y)-T_{n,q})\cap C)|\geq|C\setminus\varphi(T(y)-T_{n,q})|-|C\setminus(\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q}))|\geq 2n+2-d,$
and equality holds only if $q=0$ and $\delta_{n}\in\overline{\varphi}(T(y))$. Since $|D_{n,q}|\leq n$ and $d<3$, by the Pigeonhole Principle with $D_{n,q}$ many pigeonholes, there are at least (3d)(3-d) distinguished colors $\beta\in\Gamma^{q}_{r}\cap C$ such that $\beta\notin\varphi(T(y)-T_{n,q})$ with $\delta_{r}\in D_{n,q}\cap\overline{\varphi}(T(y))$ (different $\beta$ can have different $\delta_{r}$). Together with $|C\setminus(\Gamma^{q}\cup D_{n,q}\cup\varphi(T(y)-T_{n,q}))|=d$, the “furthermore” part holds.
Finally we prove the “consequently” part of the current claim. Since there are only two edges in $T(y_{k+2})-T(y_{k})$, clearly it holds by the “furthermore” part if $k\geq 1$. So we assume $k=0$, that is, $y=y_{0}$ being the last vertex of $T_{n,q}$. Note that if $q\geq 1$, we have $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+3$. Since there are only two edges in $T(y_{2})-T_{n,q}$, $D_{n,q}\notin\overline{\varphi}(T_{n,q})$, and $|\Gamma^{q}|\leq 2n$, we have a color $\beta\in\overline{\varphi}(T_{n,q})$ with $\beta\notin\varphi(T(y_{2})-T_{n,q})$ such that $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$, as desired. So $q=0$, and we have $|\overline{\varphi}(T_{n,q})\setminus\{\gamma_{n}\}|\geq 2n+2$ in this case. However, in this case the first edge on $T(y_{2})-T_{n,q}$ is colored by $\delta_{n}\notin\overline{\varphi}(T_{n})$. So again, since there are only two edges in $T(y_{2})-T_{n,q}$, $D_{n,q}\notin\overline{\varphi}(T_{n,q})$, and $|\Gamma^{q}|\leq 2n$, we also have a desired color $\beta\in\overline{\varphi}(T_{n,q})$ with $\beta\notin\varphi(T(y_{2})-T_{n,q})$ such that $\beta\notin\Gamma^{q}\cup D_{n,q}\cup\{\gamma_{n}\}$. ∎